OpenStudy (anonymous):

x^6 + y^6 = 1 What does: y'' = My book says y'' = (-5x^4) / (y^11) But I don't understand how. My work so far: y' = 6x^(5) + 6y^(5) * dy/dx = 0 y' = (-x^5)/(y^5) Now I use quotient rule. (f'g - fg')/(g^2) y'' = [(-5x^4)(y^5) + (x^5)(5y^4)(dy/dx)] / (y^10) Then i substituted (-x^5)/(y^5) for (dy/dx) Then I simplified, and factored. Ended up with (25x^10 - 5x^4) / (y^10)

OpenStudy (ranga):

\[\Large y" = \frac{ [(-5x^{4}y^{5} + x^{5}(5y^{4})(-x^{5}/y^{5})]}{ y{10} }\] Multiply top and bottom by y so the y^5 at the bottom in the second part of the numerator will cancel out. Also factor out -5x^4 in the numerator:

OpenStudy (ranga):

\[\Large y'' = \frac{ -5x^{4}(y^{6} + x^{6}) }{ y^{11} }\]

OpenStudy (anonymous):

can we do the product rule of -x^5/y^5

OpenStudy (ranga):

But x^6 + y^6 = 1 So y'' = -5x^4 / y^11

OpenStudy (anonymous):

ohhhhhhh gotcha but i can do the quotient rule right haha

OpenStudy (anonymous):

cant

OpenStudy (ranga):

You don't have to take the derivative anymore. You are done with the derivative. All you have to do is to simplify. I just took what you had for the second derivative and I just did the simplification part.

OpenStudy (ranga):

You had these two lines: y'' = [(-5x^4)(y^5) + (x^5)(5y^4)(dy/dx)] / (y^10) Then i substituted (-x^5)/(y^5) for (dy/dx) I did the substitution and simplified.

OpenStudy (anonymous):

okay thank you !:)

OpenStudy (ranga):

you are welcome. Glad to see you did most of the work and just wanted some clarification for the last step.

OpenStudy (anonymous):

wait i dont understand the multiplying the y to the numerator part

OpenStudy (ranga):

Let us start with the second part of the numerator: x^5(5y^4)(−x^5/y^5

OpenStudy (ranga):

Simplify it. -5x^10y^4/y^5 = -5x^10/y

OpenStudy (anonymous):

oohhhhhhhhhh haha wow thank youuuu

OpenStudy (ranga):

np. yw.

OpenStudy (anonymous):

are you a math teacher haha

OpenStudy (ranga):

not really. Software Engineer.

OpenStudy (anonymous):

nice my dads an ee haha

OpenStudy (ranga):

oh cool. He will be able to easily help you out.

OpenStudy (anonymous):

yeah exactly but hes sleeping cause he has to work and im always stuck on the same problem for hours

OpenStudy (ranga):

yeah work keeps people busy and tired when they return home. But it looks like you will do very well in calc. You did most of the work yourself and correctly.

OpenStudy (ranga):

Do well in your test. c ya.

OpenStudy (anonymous):

ok one more question i factored it out and it was -5x^4(y^5+x^6)

OpenStudy (ranga):

(-5x^4)(y^5) -5x^10/y = -5x^4(y^5 + x^6/y) = -5x^4(y^6 + x^6) / y

OpenStudy (anonymous):

and then you just multiply it by y/y?

OpenStudy (ranga):

How do you add y^5 + x^6/y ? Multiply each term by y and then divide by y. (y^6 + x^6) / y

OpenStudy (ranga):

Or if you prefer: y^5 + x^6/y Multiply the first term by y/y: y^6/y + x^6/y and you can make y the common denominator: (y^6 + x^6) / y. That y will drop down to the denominator which already has a y^10 make it a y^11.

OpenStudy (anonymous):

thank you. this helped me alot

OpenStudy (ranga):

you are welcome.